OpenStudy (anonymous):

given the arithmetic series 34+31+28+25+22+..... Find the sum of the first n terms when n=32 ?

OpenStudy (amistre64):

it appears that they are sequenced by -3 from the one before

OpenStudy (amistre64):

what do you know about the formula for an arith.seq.?

OpenStudy (anonymous):

well yea i figured d=-3 term1 was 34 and n=32 and the is sn=(t1+tn) n/2 but you need the last term to get it so you try tn=t1+ d(n-1) but i dont know how to work them out right d= common differance= -3 n= number of terms= 32

OpenStudy (amistre64):

looks about right; let me try then an = 34 -3(n-1) a32 = 34 -3(31) a32 = 34-93 a32 = -59 is what i get

OpenStudy (amistre64):

thats just for the a32 to

OpenStudy (amistre64):

to sum up; we can add the first and the last terms and multiply by how many terms/2

OpenStudy (amistre64):

there is some ploy about the numbers being seperated by more than 1 tho

OpenStudy (amistre64):

4+8+12+16 = 40 4+16 = 20*4/2 = 40.... then again, maybe not

OpenStudy (anonymous):

um? even you can't solve the math problem i posted 0.0?

OpenStudy (amistre64):

lol .... i tried to go to it and my browser froze, its on my queue tho

OpenStudy (anonymous):

oh...o.k..sorry to distrub ya!

OpenStudy (anonymous):

thank you so freiken much now i have to figure out how to do the sn=(t1+tn) n/2

OpenStudy (amistre64):

34 + 31 + ... + -56 + -59 -59 -56 + ... + 31 + 34 ------------------------- -25 -25 - ... -25 -25 looks to be -25(32)/2

OpenStudy (amistre64):

when in doubt, flip it out and do it the old school

OpenStudy (amistre64):

it tells us that we are adding up -25, well, 32 times; but that is 2 times to many so half it

OpenStudy (anonymous):

thanks man this is my first time ever on here and you made it worth it i appreciate it

OpenStudy (amistre64):

youre welcome, and good luck :)

OpenStudy (anonymous):

given the arithmetic series 34+31+28+25+22+..... find N for Sn=-12 does anybody have a clue???? do you understand this?? i wouldnt even know what equation to use? please help me

OpenStudy (anonymous):

given the arithmetic series 34+31+28+25+22+..... find N for Sn=-12 does anybody have a clue???? do you understand this?? i wouldnt even know what equation to use? please help me

OpenStudy (amistre64):

34 + x = -12 x = -46 -46 = 34 -3(n-1) -46-34 = -3(n-1) -80/-3 = n-1 80/3 + 1 = n might not be what im thinking of tho

OpenStudy (anonymous):

i know the answer i need to know how i got it like the right equation? the answer is 24

OpenStudy (anonymous):

24 is the number of terms it takes to make the sum of the sequence -12 but i need to know how to do that does anybody have a clue?

OpenStudy (amistre64):

yeah, your spose to solve for n in the equation

OpenStudy (anonymous):

do you know the equation i need to know how to solve for n

OpenStudy (amistre64):

-12 = 34 -3(n-1) -12-34 = -3(n-1) -46/-3 = n-1 15 .... it aint working out to a good number

OpenStudy (amistre64):

lets see if its 24 for starts a24 = 34 -3(23) a24 = 34 - 69 a24 = -25 -12 is NOT the 24the term

OpenStudy (amistre64):

are you using a different sequence now?

OpenStudy (amistre64):

now the SUM of the nth terms is -12

OpenStudy (anonymous):

yea i know i dont get it my teacher is crazy no i dont have any more i just need to find n if the sum = -12 i have no clue what to do but by the way thank you for being patient with me and trying

OpenStudy (amistre64):

-12 = (f+l)n/2 -24 = (34+l )n ill never learn if i dont try :) we can brunt it out if its small enough

OpenStudy (amistre64):

34 + (34-3(n-1)) n ----------------- = -12 is what we are gong for 2

OpenStudy (anonymous):

i did already thats how i came up with 24? lol i thought that equation was for geometric series????

OpenStudy (amistre64):

34n + (34-3n-3)n = -24 34n + 31n -3n^2 = -24 0 = 3n^2 -65n -24, this might help if i stuck the the math right

OpenStudy (amistre64):

nah, this equation states its arith seq.

OpenStudy (amistre64):

"given the arithmetic series 34+31+28+25+.."

OpenStudy (anonymous):

umm????? i am lost

OpenStudy (amistre64):

i already see a mistake :) 71n -3n^2 = -24

OpenStudy (amistre64):

lets start at the begining: we know that if this is going to work, then the formula for the sum of a sequance will have to fit. \[\frac{n}{2}(a1 + (a1+d(n-1)))=-12\]

OpenStudy (amistre64):

we dont know what the "last" term is, but it will have to fit into that set up

OpenStudy (anonymous):

yes??

OpenStudy (amistre64):

\[\frac{n}{2}(a1 + (a1+d(n-1)))=-12\] \[\frac{n(2)}{2}(a1 + (a1+d(n-1)))=-12(2)\] \[n(a1 + (a1+dn-d))=-24\] \[a1n + n(a1+dn-d)=-24\] \[a1n + a1n+dn^2-dn=-24\] \[2a1n+dn^2-dn=-24\] you with it so far?

OpenStudy (anonymous):

gettin there

OpenStudy (amistre64):

\[2a1n+dn^2-dn=-24\]fill in the terms we know now \[2(34)n+(-3)n^2-(-3)n=-24\] \[68n-3n^2+3n=-24\] \[71n-3n^2=-24\] \[0=3n^2-71n-24\]

OpenStudy (amistre64):

now our options for "n" will be whatever we get from the quadratic formula; disregarding any negative results

OpenStudy (amistre64):

im not sure how much you are aware of the quad formula but here it is: \[\frac{-b+\sqrt{b^2-4ac}}{2a};\text{ given }ax^2+bx+c=0\]

OpenStudy (anonymous):

wow! never seen that one before lol

OpenStudy (anonymous):

my teacher never went over this so on my test tommorow i have to do this??

OpenStudy (amistre64):

\[0=3n^2-71n-24\] \[n=\frac{71\pm\sqrt{71^2-4(3)(-24)}}{2(3)}\] \[n=\frac{71\pm\sqrt{5041-12(-24)}}{6}\] \[n=\frac{71\pm\sqrt{5041+288}}{6}\] \[n=\frac{71\pm\sqrt{5329}}{6}\] \[n=\frac{71\pm73}{6}\] \[n=\frac{-2}{6},\frac{144}{6}\] \[n=\cancel{\frac{-2}{6}},24\]

OpenStudy (amistre64):

you have to if your gonna find the n yes

OpenStudy (amistre64):

the other option is to try and err till you find something close and narrow it in

OpenStudy (amistre64):

ummmm ...... i see something that i didnt before

OpenStudy (anonymous):

so n=4 for the sum to be -12?

OpenStudy (amistre64):

24 ... but look remember up there how i got to this step? \[n(a1 + (a1+dn-d))=-24\] notice how |-24| = 24 ?

OpenStudy (amistre64):

and how we want an answer of 24?

OpenStudy (anonymous):

yea! ok im starting to click a little

OpenStudy (amistre64):

34+31+28+25 \[\frac{4}{2}(34+25)=118\] i think it was just a coincidence since 118(2) is NOT equal to 4

OpenStudy (anonymous):

you just lost me again so is there a shorter way to show my work on this problem from the beggening?

OpenStudy (amistre64):

\[\frac{n}{2}(34+[34-3(n-1)])=-12\] we would still have to find n

OpenStudy (amistre64):

we could do trial and error; start at 10 and work it out if need be. \[\frac{n}{2}(34+[34-3(n-1)])=-12\] \[n(34+34-3n+3)=-24\] \[n(71-3n)=-24\] \[10(71-3(10))=-24\] \[10(71-30)=-24\] \[10(41)=-24;not.it\]------------------------ \[20(71-3(20))=-24\] \[20(11)=-24;not.it\]------------------------ \[30(71-3(30))=-24\] \[30(-19)=-24\] \[-570=-24\] well, its between 20 and 30

OpenStudy (anonymous):

damn so what is n 24? was that right?

OpenStudy (amistre64):

\[n(71-3n)=-24\] \[25(71-3(25))=-24\] \[25(-4)=-24\] \[-100=-24\]---------------- \[24(71-3(24))=-24\] \[24(71-72)=-24\] \[24(-1)=-24\] and we find it thru trial and error

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

does the trial and error make sense?

OpenStudy (anonymous):

no treally to me im a junior in high school no one ever tought me this she just said get n lol

OpenStudy (amistre64):

well, i dont know how to make it any simpler than the trial and error.

OpenStudy (amistre64):

we take what we know; the sum formula for arithmetic seq; and fill test out "n"s till we find one that sticks in a systematic manner

OpenStudy (amistre64):

eventually it narrow down to the right choice

OpenStudy (amistre64):

anywhos; n=24 a24 = 34 -3(23) a24 = 34 -69 a24 = -35

OpenStudy (anonymous):

thank you man i would be lost with out ya

OpenStudy (amistre64):

youre welcome; it takes practice, and not really something you can just acquire over night :)

OpenStudy (anonymous):

well the test is tommorow so i can acquire as much as i can lol

OpenStudy (anonymous):